# Medical-GAT: Cancer Document Classification Leveraging Graph-Based Residual Network for Scenarios with Limited Data

### v3

@@ Abstract
Accurate classification of cancer-related medical abstracts is crucial for healthcare management and research. However, obtaining large, labeled datasets in the medical domain is challenging due to privacy concerns and the complexity of detailed clinical data. This scarcity of annotated data impedes the development of effective machine learning models for cancer document classification. To address this challenge, we present a curated dataset of 1,874 biomedical abstracts, categorized into thyroid cancer, colon cancer, lung cancer, and generic topics. Our research focuses on leveraging this dataset to improve classification performance, particularly in data-scarce scenarios. We introduce a Residual Graph Attention Network (R-GAT) with multiple graph attention layers that effectively capture the semantic information and structural relationships within lengthy cancer-related documents. Our R-GAT model is compared with various techniques, including transformer-based models such as Bidirectional Encoder Representations from Transformers (BERT) and the Robustly Optimized BERT Pretraining Approach (RoBERTa), as well as domain-specific transformer models, i.e., Bidirectional Encoder Representations from Transformers for Biomedical Text (BioBERT) and Bio+ClinicalBERT. We also evaluated deep learning-based models (e.g., CNNs, LSTMs) and traditional machine learning models (e.g., Logistic Regression, SVM) for a comprehensive comparison. Additionally, we explore ensemble approaches that combine multiple deep learning models to further enhance classification performance. Various feature extraction methods are assessed, including Term Frequency-Inverse Document Frequency (TF-IDF) with unigrams and bigrams, Word2Vec, and tokenizers from BERT and RoBERTa. The R-GAT model outperforms other techniques, achieving precision, recall, and F1 scores of 0.99, 0.97, and 0.98 for thyroid cancer; 0.96, 0.94, and 0.95 for colon cancer; 0.96, 0.99, and 0.97 for lung cancer; and 0.95, 0.96, and 0.95 for generic topics. Furthermore, the R-GAT model demonstrates better generalizability compared to both machine learning and transformer-based models. The dataset is publicly available at [^1] https://github.com/eliashossain001/MedicalAbstracts.
@@ 1 Introduction
  • We publicly release a specialized dataset of $\approx$ 1,874 cancer-related publications, categorized into thyroid, colon, and lung cancers, which can serve as a key resource for targeted healthcare research.
  • We introduced an enhanced residual graph attention network (R-GAT) to improve feature extraction and reduce information loss in complex cancer records.
  • We assessed various ML and DL models, including ensemble and transformer-based approaches, identifying their limitations and areas for improvement.
@@ 4.2.1 Classification Report of Traditional Machine Learning Models
Table 1: Classification Results for Various ML Algorithms Using Different Feature Extraction Methods
Algorithm Feature Extraction Thyroid Colon Lung Generic
@@
Decision Tree [17] TF-IDF (Unigram) 0.99 0.98 0.98 0.94 0.96 0.95 0.94 0.97 0.95 0.97 0.93 0.95
Random Forest [18] TF-IDF (Unigram) 0.99 1.00 0.99 1.00 0.98 0.99 0.99 1.00 0.99 0.99 0.99 0.99
k-nearest Neighbors [19] TF-IDF (Unigram) 1.00 0.15 0.27 0.58 0.66 0.62 0.95 0.19 0.31 0.39 0.99 0.56
Multinomial Naïve Bayes [20] TF-IDF (Unigram) 0.95 0.98 0.96 0.90 0.93 0.91 0.92 0.91 0.91 0.98 0.92 0.91
Gradient Boosting [21] TF-IDF (Unigram) 0.99 0.99 0.99 0.99 0.97 0.98 0.96 0.98 0.97 0.97 0.97 0.97
Adaptive Boosting [22] TF-IDF (Unigram) 1.00 0.69 0.82 0.99 0.95 0.97 0.95 0.99 0.97 0.73 0.96 0.83
Support Vector Machine [23] TF-IDF (Unigram) 1.00 0.94 0.97 0.88 0.98 0.93 0.99 0.94 0.96 0.96 0.97 0.96
XGBoost [24] TF-IDF (Unigram) 0.98 0.99 0.98 0.97 0.95 0.96 1.00 0.96 0.98 0.94 0.99 0.97
Logistic Regression [25] TF-IDF (Unigram) 1.00 1.00 1.00 0.99 0.96 0.97 0.99 0.99 0.99 0.95 0.98 0.96
Decision Tree [17] TF-IDF (Bigram) 0.97 0.99 0.98 0.95 0.94 0.95 0.96 0.96 0.96 0.91 0.90 0.91
Random Forest [18] TF-IDF (Bigram) 1.00 0.98 0.99 0.98 0.95 0.97 0.99 0.97 0.98 0.92 0.99 0.95
k-nearest Neighbors [19] TF-IDF (Bigram) 0.00 0.00 0.00 0.63 0.62 0.63 1.00 0.08 0.15 0.34 1.00 0.51
Multinomial Naïve Bayes [20] TF-IDF (Bigram) 0.97 1.00 0.98 0.94 0.96 0.95 0.98 0.96 0.97 0.99 0.96 0.97
Gradient Boosting [21] TF-IDF (Bigram) 1.00 0.98 0.99 0.98 0.91 0.94 0.98 0.95 0.96 0.86 0.97 0.91
Adaptive Boosting [22] TF-IDF (Bigram) 1.00 0.99 0.99 0.97 0.88 0.92 0.92 0.91 0.92 0.84 0.94 0.88
Support Vector Machine [23] TF-IDF (Bigram) 0.43 0.18 0.26 0.43 0.18 0.26 0.10 0.15 0.36 0.48 0.44 0.49
XGBoost [24] TF-IDF (Bi-gram) 0.37 0.37 0.37 0.37 0.37 0.37 0.51 0.51 0.51 0.48 0.47 0.52
Logistic Regression [25] TF-IDF (Bigram) 0.67 0.51 0.64 0.67 0.51 0.64 0.59 0.33 0.48 0.78 0.56 0.61
Decision Tree [17] Word2Vec 0.53 0.58 0.55 0.96 0.93 0.94 0.66 0.67 0.67 0.85 0.82 0.83
Random Forest [18] Word2Vec 0.66 0.68 0.67 0.96 0.96 0.96 0.76 0.77 0.76 0.88 0.85 0.86
k-nearest Neighbors [19] Word2Vec 0.58 0.67 0.62 0.94 0.96 0.95 0.72 0.68 0.70 0.89 0.80 0.84
Multinomial Naïve Bayes [20] Word2Vec 0.00 0.00 0.00 0.47 0.97 0.63 0.00 0.00 0.00 0.33 0.51 0.40
Gradient Boosting [21] Word2Vec 0.58 0.60 0.59 0.96 0.94 0.95 0.65 0.65 0.65 0.86 0.83 0.84
Adaptive Boosting [22] Word2Vec 0.22 0.51 0.30 0.33 0.01 0.03 0.53 0.37 0.43 0.48 0.56 0.52
Support Vector Machine [23] Word2Vec 0.56 0.34 0.42 0.96 0.93 0.94 0.54 0.75 0.63 0.78 0.79 0.79
XGBoost [24] Word2Vec 0.43 0.50 0.46 0.96 0.95 0.95 0.41 0.37 0.39 0.61 0.60 0.60
Logistic Regression [25] Word2Vec 0.42 0.50 0.46 0.91 0.97 0.94 0.60 0.18 0.28 0.49 0.69 0.57
@@ 4.2.2 Performance Analysis of Deep Learning Models
Table 2: Classification Results for Various Deep Learning Models Using Different Feature Extraction Methods
@@
CNN BERT Tokenizer 0.98 0.95 0.96 0.93 0.98 0.95 0.97 0.97 0.97 1.00 0.98 0.99
RNN BERT Tokenizer 0.33 0.40 0.36 0.50 0.18 0.27 0.21 0.27 0.24 0.30 0.34 0.32
LSTM BERT Tokenizer 0.51 0.80 0.62 0.70 0.97 0.81 0.64 0.08 0.14 0.94 0.82 0.88
BI-LSTM BERT Tokenizer 0.93 0.89 0.91 0.96 0.94 0.95 0.89 0.94 0.92 0.97 0.98 0.97
GRU BERT Tokenizer 0.85 0.47 0.61 0.76 0.80 0.78 0.56 0.85 0.67 0.99 0.92 0.95
Stacked LSTM BERT Tokenizer 0.00 0.00 0.00 0.00 0.00 0.00 0.24 1.00 0.38 1.00 0.01 0.02
Stacked BI-LSTM BERT Tokenizer 0.91 0.91 0.91 0.94 0.94 0.94 0.95 0.88 0.91 0.90 0.97 0.93
@@ 4.2.5 Comparative Review of Existing Studies
Proposed R-GAT Model Thyroid, Colon and Lung Biomedical Abstracts

### v4

@@ Abstract
Accurate classification of cancer-related medical abstracts is crucial for healthcare management and research. However, obtaining large, labeled datasets in the medical domain is challenging due to privacy concerns and the complexity of detailed clinical data. This scarcity of annotated data impedes the development of effective machine learning models for cancer document classification. To address this challenge, we present a curated dataset of 1,874 biomedical abstracts, categorized into thyroid cancer, colon cancer, lung cancer, and generic topics. Our research focuses on leveraging this dataset to improve classification performance, particularly in data-scarce scenarios. We introduce a Residual Graph Attention Network (R-GAT) with multiple graph attention layers that effectively capture the semantic information and structural relationships within lengthy cancer-related documents. Our R-GAT model is compared with various techniques, including transformer-based models such as Bidirectional Encoder Representations from Transformers (BERT) and the Robustly Optimized BERT Pretraining Approach (RoBERTa), as well as domain-specific transformer models, i.e., Bidirectional Encoder Representations from Transformers for Biomedical Text (BioBERT) and Bio+ClinicalBERT. We also evaluated deep learning-based models (e.g., CNNs, LSTMs) and traditional machine learning models (e.g., Logistic Regression, SVM) for a comprehensive comparison. Traditional models were included as they can sometimes capture specific patterns better and offer useful performance baselines. Additionally, we explore ensemble approaches that combine multiple deep learning models to further enhance classification performance. Various feature extraction methods are assessed, including Term Frequency-Inverse Document Frequency (TF-IDF) with unigrams and bigrams, Word2Vec, and tokenizers from BERT and RoBERTa. The R-GAT model outperforms other techniques, achieving precision, recall, and F1 scores of 0.99, 0.97, and 0.98 for thyroid cancer; 0.96, 0.94, and 0.95 for colon cancer; 0.96, 0.99, and 0.97 for lung cancer; and 0.95, 0.96, and 0.95 for generic topics, respectively. Furthermore, the R-GAT model demonstrates better generalizability compared to both machine learning and transformer-based models. The dataset is available on GitHub [^1] https://github.com/eliashossain001/MedicalAbstracts and is publicly accessible for research purposes.
@@ 1 Introduction
  • A specialized dataset of $\approx$ 1,874 cancer-related publications has been publicly released, categorized into thyroid, colon, and lung cancers, serving as a key resource for targeted healthcare research.
  • An enhanced residual graph attention network (R-GAT) has been introduced to improve feature extraction and reduce information loss in complex cancer records.
  • Various ML and DL models, including ensemble and transformer-based approaches, have been assessed, with their limitations and areas for improvement identified.
@@ 4.2.1 Classification Report of Traditional Machine Learning Models
Table 1: Classification Results for Various ML Algorithms Using Different Feature Extraction Methods. P: Precision, R: Recall, F1: F1 Score.
Model Feature Extraction Thyroid Colon Lung Generic
@@
Decision Tree TF-IDF (Unigram) 0.99 0.98 0.98 0.94 0.96 0.95 0.94 0.97 0.95 0.97 0.93 0.95
TF-IDF (Bigram) 0.97 0.99 0.98 0.95 0.94 0.95 0.96 0.96 0.96 0.91 0.90 0.91
Word2Vec 0.53 0.58 0.55 0.96 0.93 0.94 0.66 0.67 0.67 0.85 0.82 0.83
Random Forest TF-IDF (Unigram) 0.99 1.00 0.99 1.00 0.98 0.99 0.99 1.00 0.99 0.99 0.99 0.99
TF-IDF (Bigram) 1.00 0.98 0.99 0.98 0.95 0.97 0.99 0.97 0.98 0.92 0.99 0.95
Word2Vec 0.66 0.68 0.67 0.96 0.96 0.96 0.76 0.77 0.76 0.88 0.85 0.86
K-NN TF-IDF (Unigram) 1.00 0.15 0.27 0.58 0.66 0.62 0.95 0.19 0.31 0.39 0.99 0.56
TF-IDF (Bigram) 0.00 0.00 0.00 0.63 0.62 0.63 1.00 0.08 0.15 0.34 1.00 0.51
Word2Vec 0.58 0.67 0.62 0.94 0.96 0.95 0.72 0.68 0.70 0.89 0.80 0.84
Multinomial NB TF-IDF (Unigram) 0.95 0.98 0.96 0.90 0.93 0.91 0.92 0.91 0.91 0.98 0.92 0.91
TF-IDF (Bigram) 0.97 1.00 0.98 0.94 0.96 0.95 0.98 0.96 0.97 0.99 0.96 0.97
Word2Vec 0.00 0.00 0.00 0.47 0.97 0.63 0.00 0.00 0.00 0.33 0.51 0.40
Gradient Boosting TF-IDF (Unigram) 0.99 0.99 0.99 0.99 0.97 0.98 0.96 0.98 0.97 0.97 0.97 0.97
TF-IDF (Bigram) 1.00 0.98 0.99 0.98 0.91 0.94 0.98 0.95 0.96 0.86 0.97 0.91
Word2Vec 0.58 0.60 0.59 0.96 0.94 0.95 0.65 0.65 0.65 0.86 0.83 0.84
AdaBoost TF-IDF (Unigram) 1.00 0.69 0.82 0.99 0.95 0.97 0.95 0.99 0.97 0.73 0.96 0.83
TF-IDF (Bigram) 1.00 0.99 0.99 0.97 0.88 0.92 0.92 0.91 0.92 0.84 0.94 0.88
Word2Vec 0.22 0.51 0.30 0.33 0.01 0.03 0.53 0.37 0.43 0.48 0.56 0.52
SVM TF-IDF (Unigram) 1.00 0.94 0.97 0.88 0.98 0.93 0.99 0.94 0.96 0.96 0.97 0.96
TF-IDF (Bigram) 0.43 0.18 0.26 0.43 0.18 0.26 0.10 0.15 0.36 0.48 0.44 0.49
Word2Vec 0.56 0.34 0.42 0.96 0.93 0.94 0.54 0.75 0.63 0.78 0.79 0.79
XGBoost TF-IDF (Unigram) 0.98 0.99 0.98 0.97 0.95 0.96 1.00 0.96 0.98 0.94 0.99 0.97
TF-IDF (Bigram) 0.37 0.37 0.37 0.37 0.37 0.37 0.51 0.51 0.51 0.48 0.47 0.52
Word2Vec 0.43 0.50 0.46 0.96 0.95 0.95 0.41 0.37 0.39 0.61 0.60 0.60
Logistic Regression TF-IDF (Unigram) 1.00 1.00 1.00 0.99 0.96 0.97 0.99 0.99 0.99 0.95 0.98 0.96
TF-IDF (Bigram) 0.67 0.51 0.64 0.67 0.51 0.64 0.59 0.33 0.48 0.78 0.56 0.61
Word2Vec 0.42 0.50 0.46 0.91 0.97 0.94 0.60 0.18 0.28 0.49 0.69 0.57
@@ 4.2.2 Performance Analysis of Deep Learning Models
Table 2: Classification Results for Various Deep Learning Models Using Different Feature Extraction Methods. P: Precision, R: Recall, F1: F1 Score.
@@
BERT Tokenizer 0.91 0.91 0.91 0.94 0.94 0.94 0.95 0.88 0.91 0.90 0.97 0.93
@@ 4.2.5 Comparative Review of Existing Studies
R-GAT Model Thyroid, Colon and Lung Biomedical Abstracts